Your work must be prepared in an excel format. Please bring a hard copy of your work to class. If you will not be attending class, submit your work via email. Late submission will be subject to a one-time penalty of 20 percent of the grade.

You are evaluating two different silicon wafer milling machines. The Techron I costs $265,000, has a 3-year life, and has pretax operating costs of $41,000 per year. The Techron II costs $330,000, has a 5-year life, and has pretax operating costs of $52,000 per year. For both milling machines, we use straight line depreciation to zero over the project’s life and assume a salvage value of $25,000. If your tax rate is 21 percent and your discount rate is 9 percent, compute the EAC for both machines. Which machine do you prefer? Why?
Answer #1

a). For Techron I;

Depreciation = Cost of the System / Useful life years = $265,000 / 3 = $88,333.33

OCF = [-Costs * (1 - t)] + [Depreciation * t]

= [-$41,000 * (1 - 0.21)] + [$88,333.33 * 0.21] = -$32,390 + $18,550 = -$13,840

NPV = PV of Cash Inflows - PV of Cash Outflows

PV of Cash Outflows = Annual OCF * [{1 - (1 + r)-n} / r] + Cost

= $13,840 * [{1 - (1 + 0.09)-3} / 0.09] + $265,000

= $13,840 * [0.2278 / 0.09] + $265,000

= [$13,840 * 2.5313] + $265,000 = $35,033.12 + $265,000 = $300,033.12

NPV = $0 - $300,033.12 = -$300,033.12

EAC = NPV / [{1 - (1 + r)-n} / r]

= -$300,033.12 / [{1 - (1 + 0.09)-3} / 0.09]

= -$300,033.12 / [0.2278 / 0.09] = -$300,033.12 / 2.5313 = -$118,529.51

For Techron II;

Depreciation = Cost of the System / Useful life years = $330,000 / 5 = $66,000

OCF = [-Costs * (1 - t)] + [Depreciation * t]

= [-$52,000 * (1 - 0.21)] + [$66,000 * 0.21] = -$41,080 + $13,860 = -$27,220

NPV = PV of Cash Inflows - PV of Cash Outflows

PV of Cash Outflows = Annual OCF * [{1 - (1 + r)-n} / r] + Cost

= $27,220 * [{1 - (1 + 0.09)-5} / 0.09] + $330,000

= $27,220 * [0.3501 / 0.09] + $330,000

= [$27,220 * 3.8897] + $330,000 = $105,876.31 + $330,000 = $435,876.31

NPV = $0 - $435,876.31 = -$435,876.31

EAC = NPV / [{1 - (1 + r)-n} / r]

= -$435,876.31 / [{1 - (1 + 0.09)-5} / 0.09]

= -$435,876.31 / [0.3501 / 0.09] = -$435,876.31 / 3.8897 = -$112,060.51

If the machine will be continually replaced, we should choose Techron II since it has the more positive EAC.